OpenStudy (anonymous):

simplify the following expresssion. √-11√-55

sam (.sam.):

Just multiply\[ \sqrt{ (-11)(-55)}\]

OpenStudy (callisto):

\[\sqrt{-11}\sqrt{-55} = \sqrt{-11\times -55} = \sqrt{11^2\times 5} = \sqrt{11^2} \sqrt{5} =?\]

OpenStudy (anonymous):

11√5

OpenStudy (callisto):

Wow... thank you for giving the answer to the question... But it's better to leave some work for the asker, so that he/she can really to something for the question!

OpenStudy (anonymous):

got you